25. Today (T=0), an investor purchased a 20 year bond with a 5.00% coupon and a face value of $100,000 for $106,550. In six months (T=0.5) interest rates have decreased by 0.50% and the investor decides to sell the bond immediately after receiving the first coupon payment. What is the investor’s total gain (loss) on the bond? HINT: Total Gain (Loss) = Price Change in Bond + Coupon

  1. ($6,548)

  2. ($6,048)

  3. $7,130

  4. $7,602

  5. $7,630
Answer #1
A1 B C D E F G H I J K
2
3 Calculation of yield to maturity:
4 Face value $100,000
5 Coupon rate 5%
6 Current Price $106,550
7 Maturity 20 years
8 Semi-Annual Coupon $2,500
9 Semi-annual Period 40
10
11 Cash flow to investor will be as follows:
12 Semi-Annual Period 0 1 2 3 4 40
13 Cash flow ($106,550.00) $2,500 $2,500 $2,500 $2,500 $2,500 $102,500
14
15 Yield to maturity is the rate at which if future NPV to Investor will be zero.
16
17 Rate(nper,pmt,PV, [fv],type) function of excel can be used to find the yield to maturity as follows:
18 nper 40
19 pmt $2,500
20 PV ($106,550.00)
21 FV $100,000
22
23 Semi annual Yield to maturity 2.25% =RATE(D18,D19,D20,D21)
24 Yield to maturity =2* Semi-annual Yield to maturity
25 4.50% =D33*2
26
27 Hence current interest rate is 4.50%
28
29
30
31 Calculation of Price of the Bond after 1st Payment:
32 Par value (F) $100,000
33 Semiannual Coupon Rate 5%
34 Market demanded return (Yield to maturity) 4.00% =D27-0.5%
35 Original term of Bond 20 Years
36 Number of coupons Paid 1
37 Interest is paid twice a year i.e. semiannual.
38 Semiannual coupon (C) $2,500.00
39 Remaining Semiannual Period (n) 39
40 Semiannual YTM (i) 2.00%
41 Current Value of the bond can be calculated by finding the present value of cash flows of bonds.
42 Cash Flow of Bonds can be written as follows:
43 Semiannual Period 0 1 2 3 4 39
44 Cash Flow of Bonds $2,500.00 $2,500.00 $2,500.00 $2,500.00 $2,500.00 $102,500.00
45
46 Value of Bond =C*(P/A,i,n)+F*(P/F,i,n)
47 Where, C is Semiannual coupon, F is par value of bond, i is semiannual market rate and n is total semiannual periods.
48
49 Value of Bond after first coupon payment =C*(P/A,i,n)+F*(P/F,i,n)
50 =2500*(P/A,2%,39)+100000*(P/F,2%,39)
51 $113,453.03 =D38*PV(D40,D39,-1,0)+D32*(1/((1+D40)^D39))
52 Hence Selling price of bond is $113,453.03
53
54 Calculation of Total Gain:
55 Purchase Price of Bond $106,550
56 Selling Price of Bond $113,453.03
57 Coupon Received $2,500.00
58
59 Total Gain =(Selling Price - Purchase Price) + Coupon Received
60 $9,403.03 =(D56-D55)+D57
61
62 Hence total gain is $9,403.03
63 None of the answers is matching.
64 Please check the options given again.
65